OpenStudy (anonymous):

divide and simp 5 lil 19 5 lil 4 type type an exponential notation with positive exponents

OpenStudy (anonymous):

Is "5 lil 19" 5^19? \[\Large \frac{{{x^n}}}{{{x^m}}} = {x^{n - m}}\]So\[\Large \frac{{{5^{19}}}}{{{5^4}}} = {5^{19 - 4}} = {5^{15}}\]

OpenStudy (anonymous):

thanks good buddy

OpenStudy (anonymous):

No problemo.